OpenStudy (anonymous):

Can you walk me through this word problem? You order baseball caps with your team logo. Each cap costs $9.95. There is a design fee of $60.00 and an embroidery charge of $1.50 per cap. Write an expression to find the total cost of "c" baseball caps. What is the total cost for 20 baseball caps?

OpenStudy (chaise):

Cost = 60 + n(9.95+1.50) Where n = caps you want. Cost = 60 + 20(11.45) Cost = 60 + 229 Cost = $289 Does this make sense? I hope I did this right. Instead of 9.95 + 1.50 you can change this to 11.45, I just wanted to make it explicit how I got 11.45, rather than just writing it. Seeing as though you asked HOW to do it, rather than just an answer.

OpenStudy (chaise):

So you're paying for the design fee once, as once it's designed it cannot be designed again. You're paying for each cape (9.95), plus the design to be put onto the cap, (1.50)

OpenStudy (lalaly):

i agree with chaise:)

OpenStudy (anonymous):

So the 60+11.45 is going to be there for other amounts of caps as well?